Thinfilms morning report

james haydon jhaydon at snf.stanford.edu
Tue Jan 3 16:31:31 PST 2006


           

Thinfilms Morning Report 1/3/2006

 

 

Gryphon:            bad turbo or controller

Innotec;            rebuilt lift, centered carousel, changed shields, 
installed stop to locate dome

STS Dep:            ok

Implanter:            ok to use but still has some problem with noise on 
the HV.

Metalica:            ok

SCT:                ok

CPD:                ok                   

WbGen-CTB:            ok

WbGen-RHP:            ok

WbGen:-LHP:            waiting for parts to rebuild hot plate

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            replaced RTD on left hot pot

WbDif:            replaced RTD on left hot pot

WbNonmetal:            replaced both RTD's

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok

 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/thinfilms/attachments/20060103/60cb2faf/attachment.html>


More information about the thinfilms mailing list